# Misty: UI Prototyping Through Interactive Conceptual Blending

### v2

@@ 8. Conclusion
We built Misty, a system that embodies the conceptual blending and enables designers and developers to blend concepts from UI examples into their own work. We operationalized conceptual blending within Misty through a multi-modal AI model—postulating that these models have some capability to to interpret and apply conceptual blending. Although imperfect and occasionally uneven in its generation, these models nevertheless demonstrated a remarkable ability to apply blends and even produce surprising and serendipitous results. Participants in our exploratory first-use study were encouraged by Misty’s capabilities. Creativity is a kaleidoscope, and Misty demonstrates the potential for tools that blur the boundaries between developers and designers.

### v3

@@ 9. Conclusion
We built Misty, a system that embodies the conceptual blending and enables designers and developers to blend concepts from UI examples into their own work. We operationalized conceptual blending within Misty through a multi-modal AI model—postulating that these models have some capability to interpret and apply conceptual blending. Although imperfect and occasionally uneven in its generation, these models nevertheless demonstrated a remarkable ability to apply blends and even produce surprising and serendipitous results. Participants in our exploratory first-use study were encouraged by Misty’s capabilities. Creativity is a kaleidoscope, and Misty demonstrates the potential for tools that blur the boundaries between developers and designers.